OpenStudy (anonymous):

Complete the solution of the equation. Find the value of y when x equals 1. 9x + 7y = -12

OpenStudy (anonymous):

just 1 instead of x the solve for y

OpenStudy (anonymous):

9x+7y=-12 9(1)+7y=-12 9+7y =-12

OpenStudy (anonymous):

9(1) + 7y = -12 -12 - 9 = -21 7y = -21 -12 / 7 = the answer... i suppose.

OpenStudy (anonymous):

@Ericaa_ ?